OpenStudy (anonymous):

Write each equation in slope intercept form: 4x-2y=7

OpenStudy (anonymous):

is in standard form. to write it in slope intercept form u need to leave in this form ( y=mx+b) so solve for y 4x-2y=7-- subtarct 4x ur left with -2y=7-4x now divide by -2 to leave y alone, so ur left with y=7/2-4/2x which is eqaul to y=-2x+7/2 ( this is the answers in slope intercept form)